(3.3.25) Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find P40.
0.1
0.2
0.3
0.5
0.6
0.6
0.7
0.7
0.8
0.8
0.8
0.8
0.8
0.9
0.9
1.3
1.4
1.5
1.6
1.6
1.6
2.1
2.1
2.4
2.4
2.5
3.4
3.6
4.1
4.2
4.5
4.6
5.7
6.7
7.7
8.1
8.5
12.6
12.6
12.8
12.8
13.3
13.3
13.7
14.2
14.3
15.1
15.5
15.6
28.6
P40=? Mbps
1
Expert's answer
2020-12-07T20:52:17-0500

The data are in ascending order.

Compute the position of the pth percentile (index "i" ):


"i=\dfrac{p}{100}\times n=\dfrac{40}{100}\times50=20"

The index "i"  is an integer ⇒ the 40th percentile is the average of the values in the 19th and 20th positions (1.6 and 1.6 respectively)

Answer: the 40th percentile is (1.6 + 1.6) / 2 = 1.6


P40=1.6 Mbps